OpenStudy (anonymous):

A line passes through (3, 5) and (6, 14). What is the equation of the line in point-slope form?

Directrix (directrix):

Use the attachment to help understand how to find the slope of the line determined by the two points. Post what you get. @muellbrockc00

OpenStudy (anonymous):

y-5=3(x-3)

Directrix (directrix):

Lemme see what I get. Hold on.

Directrix (directrix):

I agree. I was thinking that we needed to simplify but if we did, we could most likely end up with the line in slope-intercept form which is not the question. @muellbrockc00

Directrix (directrix):

If y - 14 = 3(x - 6) is an option, then we may have to reconsider.

OpenStudy (anonymous):

well it says my answer or your answer, so like there saying that it can be eaither one i guess

Directrix (directrix):

Hey, could you post all the options for a few minutes so that I could look at them. Then, you can click on delete and take them away.